Calculate the critical frequency for the E layer if maximum ionic densities are 3 x 106

A.15.58 x 106

B.15.58 kHz

C.1.58 x 103

D.None of the above


1
Expert's answer
2021-01-19T07:10:31-0500

The critical frequency for the E layer can be calculated as follows:


"f_c=9\sqrt{N_{max}},"

"f_c=9\cdot\sqrt{3\cdot10^6\ \dfrac{electrons}{m^3}}=15.58\cdot10^3\ Hz =15.58\ kHz."

Answer:

"f_c=15.58\ kHz."